# QueenV2: Future of Quantum Circuit Simulation Seven Seas of Rhye

### v1

@@ II-B Optimized Techniques for High-Performance Computing
To achieve effective state vector-based QCS, it is essential to incorporate complementary optimization techniques such as optimizing data locality, minimizing unnecessary operations, and rearranging the order of gates. he most effective mechanisms known to date include but are not limited to, gate fusion, qubit reordering, and in-cache operations.
@@ III QueenV2
As illustrated in Figure 1, the QueenV2 workflow fully inherits from the first generation while incorporating significant modifications in both the all-in-one optimization module (AIO) and the all-in-cache simulation module (AIC). As the author is currently learning about other research topics, a quick overview of the techniques is provided for peers and researchers. Once this phase is completed, any implementation details will be shared soon.
@@ III-A All-in-one circuit module
The overall schematic is illustrated in Figure 2. To ensure that everyone can implement the qubit reordering mechanism, please prepare a bi-directly corresponding hash table for all qubits. The orange and purple regions represent the gate block sections for Tensor Product Booster. The diagonal gates will be merged as much as possible when the conditions are met, and the 11th gate needs to be excluded due to it spanning across ranks. Moreover, it is evident from the diagram that in-memory swapping (IMS) is no longer necessary in current QueenV2, with its use limited to instances of cross-rank swapping (CRS).
@@ III-B All-in-cache simulation module
The mechanism for reading from and writing back device memory has been incorporated through the implementation of instructions similar to the PEDP and PEXT instruction sets. Special thanks to Wang, my dear schoolmate, for his contributions in this area. In CRS operations, IMS operations are still utilized due to the necessity of arranging sub-state vectors contiguously, facilitating efficient data transmission across multiple ranks to maximize the data transfer throughput.

### v2

@@ II-B Optimized Techniques for High-Performance Computing
To achieve effective state vector-based QCS, it is essential to incorporate complementary optimization techniques such as optimizing data locality, minimizing unnecessary operations, and rearranging the order of gates. The most effective mechanisms known to date include but are not limited to, gate fusion, qubit reordering, and in-cache operations.
@@ III QueenV2
As illustrated in Figure 1, the QueenV2 workflow fully inherits from the first generation while incorporating significant modifications in both the all-in-one optimization module (AIO) and the all-in-cache simulation module (AIC). As the author is currently learning about other research topics, a quick overview of the techniques is provided for peers and researchers. Once this phase is completed, any implementation details and knowledge corrections will be shared soon.
@@ III-A All-in-one circuit module
The overall schematic is illustrated in Figure 2. To ensure that everyone can implement the qubit reordering mechanism, please prepare a bi-directly corresponding hash table for all qubits. The orange and purple regions represent the gate block sections for Tensor Product Booster (TPB). The diagonal gates will be merged as much as possible when the conditions are met, and the 11th gate needs to be excluded due to it spanning across ranks. The 3rd gate is not within the search scope because it is located in the TPB#1 section. Moreover, it is evident from the diagram that in-memory swapping (IMS) is no longer necessary in current QueenV2, with its use limited to instances of cross-rank swapping (CRS).
@@ III-B All-in-cache simulation module
The mechanism for reading from and writing back device memory has been incorporated through the implementation of instructions similar to the PEDP and PEXT instruction sets [^2] If NVIDIA implements similar acceleration instructions, this part could benefit from further optimization.. Special thanks to Wang, my dear schoolmate, for his contributions in this area. In CRS operations, IMS operations are still utilized due to the necessity of arranging sub-state vectors contiguously, facilitating efficient data transmission across multiple ranks to maximize the data transfer throughput.